OpenStudy (anonymous):

let f(x)=x2-1 and g(x)=3x-2 (f+g)(2)

OpenStudy (dumbcow):

f+g literally means add the 2 functions together (x2 - 1) + (3x -2) = x2 + 3x - 3 then evaluate at x=2 2^2 + 3(2) - 3 4 + 6 - 3 10 - 3 7

OpenStudy (anonymous):

You can also evaluate each function individually and add them together. f(2)+g(2) = 3+4 = 7

OpenStudy (anonymous):

Oh!! That makes sense. Thank you so much!

OpenStudy (anonymous):

Thank you both! God bless<><